You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130402 (05 Apr 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 April 1 - April 2

Observer: E. Sani, M. Faccini, A. Bongiorno
Telescope Operator: S. Anderson
Support Astronomer: M. Edwards
Instrument: LUCI img - MODS

Summary

At the beginning of the night we have thin sky conditions with 50% humidity and wind average speed 10 m/s. Seeing is at 0.8/0.9
At the beginning of the scientific operations the average seeing on the GCS is 0.9.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard PG0823 dual - airmass 1.08
Fritz_04D GG190 dual 3h20min COMPLETE
Annibali_23A NGC4449 dual 1h30min the last 45min are affected by bad seeing and thick clouds passing. To be repeated.
Covino_10 CXOJ1348+26 dual 2h average seeing within the specs. Moon illumination 60% at 80deg from the target

Details

00:26 We start by performing MODS calibrations
  • execMODS grpixflats.cal mods1r 0002 - 0006 mods1b 0002- 0011 standard slitless flats for all programs. Are GOOD
  • execMODS grflats.cal mods1r 0007 - 0021 mods1b 0012- 0021 Fritz_04 Flats, counts too high especially for the VFlat.
  • mods1r 0007-0011 mods1b 0012-0016 are GOOD. We repeat the saturated flats at the end of the night.

01:20 We run some test to find the proper balance between the lamp power and exposure time:
  • mods1r 0022-0024 test for VFlat lamp (changing lamp power and exptime), with the red channel. GOOD with VFlat power 8.0 and exptime 4s
  • mods1r 0025 mods1b 0022 test QTH1 + QTH2 lamps, 0.3s red-channel, 2s blue channel. Low counts
  • mods1r 0026 mods1b 0023 test QTH1 + QTH2 lamps, 0.5s red-channel, 4s blue channel. Blue OK, low in the red.
  • mods1r 0027 test QTH1 + QTH2 lamps power 8.0, 0.5s low counts.
  • mods1r 0028 test QTH1 + QTH2 lamps power 8.0, 1.5s saturated.

01:52 The TO opens the dome

02:21 Slew to standard star for Fritz_04D
  • acqMODS PG0823.acq
  • modsAlign -r mods1r 0029
  • shell RIP, we haveto send the preset again

02:33 acqMODS PG0823.acq mods1r 0030-0032
  • modsAlign -r mods1r 0030

02:42 execMODS PG0823.obs mods1r 0033-0034 mods1b 0024
  • we are still in the twilight, we repeat the spectra integration also after the end of twilight

03:03 execMODS PG0823.obs
  • no images taken

03:04 TO asked to pause due to shell RIP of the secondary mirror

03:08 execMODS PG0823.obs mods1r 0035-0036 mods1b 0025

  • preset cancel to unwrap the rotator

03:24 acqMODS GG190.acq mods1r 0037

  • acq canceled due to a MODS error (red GO exposure already in progress)

03:30 acqMODS GG190.acq mods1r 0038-0040
  • modsAlign -l mods1r 0038 mods1r 0039

03:48 execMODS GG190.obs mods1r 0041-0044 mods1b 0026-0029
  • during this exposure the average seeing on the GCS increases from 0.85 to 1.5 (see attached plot)

07:21 acqMODS mods.4449_runa.acq mods1r 0045-0047
  • modsAlign mods.4449_runa.mms mods1r 0045 mods1r 0046

07:41 execMODS mods.4449_runa_short.obs mods1r 0048-49 mods1b 0030-31
  • we run a shortened ob to execute only the remaining integration time.
  • Thick + thin clouds are coming. The average seeing on GCS is 1.5

9:18 Slew to Covino_10 scientific target CXO J1348+26
  • acqMODS 10.acq mods1r 0050-0054
  • we get (twice) the same error as before, and re-run the acquisition.
  • at the third presetting, we get only the field image mods1r 0050 and the slit image is missing.
  • modsAlign -l mods1r 0051 mods1r 0052

09:59 execMODS 10.obs mods1r 0055-0058 mods1b 0032-0035
  • The sky conditions are better at the beginning of this scientific integration, with the clouds mostly gone.
  • The average seeing at the beginning of integration is 1.3 on the GCS.
  • The average seeing at the end of integration is 1.3 on the GCS.

12:13 The TO closes the DOME

12:23 We start MODS flats and arcs
  • execMODS grflats.cal for Fritz_04 program mods1r0059-0063 mods1b0036-0040== with QTH1 + QTH2 GOOD
  • mods1r0064-0068 with VFlat GOOD
  • execMODS grlamps.cal for Covino_10 mods1r0069-0077 mods1b0041-0049
  • execMODS bias.cal mods1r0078-0082 mods1b0050-0054

-- EleanoraSani - 01 Apr 2013
I Attachment Action Size Date Who Comment
night_par1.pngpng night_par1.png manage 210 K 02 Apr 2013 - 06:54 MarcoFaccini  
plot_sky_5.4_6.4.pngpng plot_sky_5.4_6.4.png manage 206 K 02 Apr 2013 - 06:55 MarcoFaccini  
sky_9.04.pngpng sky_9.04.png manage 192 K 02 Apr 2013 - 09:05 MarcoFaccini  
sky_conditions_10.00.pngpng sky_conditions_10.00.png manage 189 K 02 Apr 2013 - 10:03 MarcoFaccini  
Topic revision: r10 - 05 Apr 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback